# ConFIG: Towards Conflict-free Training of Physics Informed Neural Networks

### v1

@@ A.11 Training Details
The training of PINNs in the current research follows the established conventions. The neural networks are fully connected with 4 hidden layers and 50 channels per layer. The activation function is the Tanh function, and all weights are initialized with Xavier initialization [47]. Data points are sampled using Latin-hypercube sampling and updated in each iteration. The extended training run case shown in Fig. 10 uses a constant learning rate of $10^{-4}$. All other cases follow a cosine decay strategy with the initial and final learning rate of $10^{-3}$ and $10^{-4}$, respectively. We also add a learning rate warm-up of 100 epochs for each training. All the methods except M-ConFIG use Adam optimizer. The hyper-parameters of the Adam optimizer are set as $\beta_{1}$=0.9, $\beta_{2}=0.999$, and $\varepsilon=10^{-8}$, respectively. The number of data points and training epochs for each case are listed in Tab. 20.
@@
Kovasznay flow 20000 1000
Beltrami flow 25000 5000 5000
@@
Our experiments are based on the official test code of the FAMO method, and our ConFIG method implemented in the corresponding framework. For the details of the configurations, please refer to the original FAMO paper [40] and its official repository: https://github.com/Cranial-XIX/FAMO (MIT License).
Experiments compute resources.
All the experiments in this study were conducted using an NVIDIA RTX A5000 GPU with 24 GB of memory. Each PINN experiment discussed can be completed within a few hours on any modern GPU with more than 4 GB of memory. For the CelebA MTL test, a GPU with more than 12 GB of memory is required, and a single training run may take 1-2 days, depending on GPU performance.

### v2

@@ A.12 Training Details
The training of PINNs in the current research follows the established conventions. The neural networks are fully connected with 4 hidden layers and 50 channels per layer. The activation function is the tanh function, and all weights are initialized with Xavier initialization (Glorot & Bengio, 2010). Data points are sampled using Latin-hypercube sampling and updated in each iteration. The extended training run case shown in Fig. 10 uses a constant learning rate of $10^{-4}$. All other cases follow a cosine decay strategy with the initial and final learning rate of $10^{-3}$ and $10^{-4}$, respectively. We also add a learning rate warm-up of 100 epochs for each training. All the methods except M-ConFIG use the Adam optimizer. The hyper-parameters of the Adam optimizer are set as $\beta_{1}$=0.9, $\beta_{2}=0.999$, and $\varepsilon=10^{-8}$, respectively. The number of data points and training epochs for each case are listed in Tab. 23.
@@
Kovasznay flow 20000 1000 $10^{5}$
Beltrami flow 25000 5000 5000 $10^{5}$
PNd 20000 5000 $10^{5}$
HeatMS 20000 2000 2000 $10^{5}$
KS 20000 500 500 $10^{5}$
@@
Our experiments are based on the official test code of the FAMO method, and our ConFIG method implemented in the corresponding framework. For the details of the configurations, please refer to the original FAMO paper (Liu et al., 2023) and its official repository: https://github.com/Cranial-XIX/FAMO (MIT License).
Compute resources.
All the experiments in this study were conducted using an NVIDIA RTX A5000 GPU with 24 GB of memory. Each PINN experiment completes training within a few hours on a typical GPU with more than 4 GB of memory. For the CelebA MTL test, a GPU with more than 12 GB of memory is required, and a single training run takes ca. 1-2 days.
A.13 Ablation study on training hyperparameter
Although we utilize a momentum-based optimizer for the training, the learning rate may still affect the training process, especially considering that our ConFIG and M-ConFIG methods consistently change the gradient vector during the training. Thus, we perform an ablation study on different learning rates. As shown in Tab. 24 and Tab. 25, our methods consistently outperform other methods with different learning rate configurations.
Table 24: The performance of different methods with different cosine decay learning rates in Burgers two-loss test
$10^{-3}\rightarrow 10^{-4}$ $10^{-4}\rightarrow 10^{-5}$
Adam $1.484\pm 0.061$ $3.076\pm 1.201$
PCGrad $1.344\pm 0.019$ $2.223\pm 0.355$
IMTL-G $1.339\pm 0.024$ $1.688\pm 0.018$
MinMax $1.889\pm 0.143$ $3.980\pm 0.798$
ReLoBRaLo $1.419\pm 0.053$ $5.057\pm 1.821$
LRA $353.796\pm 114.972$ $819.730\pm 56.925$
ConFIG $1.308\pm 0.008$ $1.887\pm 0.283$
M-ConFIG $\mathbf{1.277\pm 0.035}$ $\mathbf{1.681\pm 0.106}$
Table 25: The performance of different methods with different constant learning rates in Burgers two-loss test
$\gamma=10^{-3}$ $\gamma=10^{-4}$ $\gamma=10^{-5}$
Adam $1.398\pm 0.021$ $3.076\pm 1.201$ $420.511\pm 95.578$
PCGrad $1.398\pm 0.018$ $2.223\pm 0.355$ $51.664\pm 7.753$
IMTL-G $1.385\pm 0.042$ $1.688\pm 0.018$ $192.151\pm 123.903$
MinMax $1.889\pm 0.120$ $3.980\pm 0.798$ $7.078\pm 1.672$
ReLoBRaLo $1.477\pm 0.061$ $5.057\pm 1.821$ $10.333\pm 1.905$
LRA $367.944\pm 98.322$ $819.730\pm 56.925$ $895.965\pm 1.671$
ConFIG $1.373\pm 0.006$ $1.887\pm 0.283$ $103.239\pm 40.287$
M-ConFIG $\mathbf{1.354\pm 0.019}$ $\mathbf{1.681\pm 0.106}$ $\mathbf{4.097\pm 1.310}$
In the current study, the points are sampled dynamically from the internal domain and boundaries during training. Thus, the number of data samples represents the training batch size. Meanwhile, each loss-specific gradient is evaluated through the data samples at the internal domain and boundaries. Thus, the number of sample points and the relative ratio between the number of points at the boundary and the internal domain may affect the quality of the gradient, further affecting the final training results. Here, we also perform an ablation study on the data samples, and the results are summarized in Fig. 26, Tab. 26 and Tab. 27. Our methods consistently outperform other methods with different configurations of data samples, showing the robustness of our methods.
Figure 26: The relative improvements of different methods with different numbers and different ratios of data samples during training
Table 26: The performance of different methods with different numbers of data samples during training
$\begin{matrix}n_{\mathcal{N}}=5000\\ n_{\mathcal{B}}=n_{\mathcal{I}}=125\end{matrix}$ $\begin{matrix}n_{\mathcal{N}}=10000\\ n_{\mathcal{B}}=n_{\mathcal{I}}=250\end{matrix}$ $\begin{matrix}n_{\mathcal{N}}=20000\\ n_{\mathcal{B}}=n_{\mathcal{I}}=500\end{matrix}$
Adam $1.494\pm 0.060$ $1.484\pm 0.061$ $1.438\pm 0.031$
PCGrad $1.350\pm 0.015$ $1.344\pm 0.019$ $1.331\pm 0.037$
IMTL-G $1.368\pm 0.012$ $1.339\pm 0.024$ $1.275\pm 0.003$
MinMax $1.943\pm 0.212$ $1.889\pm 0.143$ $1.846\pm 0.135$
ReLoBRaLo $1.478\pm 0.067$ $1.419\pm 0.053$ $1.381\pm 0.043$
LRA $340.161\pm 135.263$ $353.796\pm 114.972$ $374.574\pm 84.340$
ConFIG $1.307\pm 0.014$ $1.308\pm 0.008$ $1.295\pm 0.014$
M-ConFIG $\mathbf{1.291\pm 0.017}$ $\mathbf{1.277\pm 0.035}$ $\mathbf{1.262\pm 0.004}$
Table 27: The performance of different methods with different ratios of data samples during training
$\begin{matrix}n_{\mathcal{N}}=5000\\ n_{\mathcal{B}}=n_{\mathcal{I}}=250\end{matrix}$ $\begin{matrix}n_{\mathcal{N}}=10000\\ n_{\mathcal{B}}=n_{\mathcal{I}}=250\end{matrix}$ $\begin{matrix}n_{\mathcal{N}}=20000\\ n_{\mathcal{B}}=n_{\mathcal{I}}=250\end{matrix}$
Adam $1.470\pm 0.012$ $1.484\pm 0.061$ $1.455\pm 0.044$
PCGrad $1.358\pm 0.026$ $1.344\pm 0.019$ $1.355\pm 0.013$
IMTL-G $1.328\pm 0.036$ $1.339\pm 0.024$ $1.342\pm 0.027$
MinMax $1.963\pm 0.133$ $1.889\pm 0.143$ $1.843\pm 0.152$
ReLoBRaLo $1.469\pm 0.014$ $1.419\pm 0.053$ $1.393\pm 0.052$
LRA $362.388\pm 110.455$ $353.796\pm 114.972$ $350.266\pm 127.639$
ConFIG $1.298\pm 0.014$ $1.308\pm 0.008$ $1.300\pm 0.010$
M-ConFIG $\mathbf{1.289\pm 0.015}$ $\mathbf{1.277\pm 0.035}$ $\mathbf{1.265\pm 0.026}$